OpenStudy (anonymous):

y = log x If y = 10, then what is x?

OpenStudy (anonymous):

is that log to the base 10?

OpenStudy (anonymous):

if so then x = 10^10

OpenStudy (anonymous):

what they are asking is for 10=log(x) what does x equal. Do you know how to get rid of the log so you just have x on one side?

OpenStudy (anonymous):

so it'll give us y = log x 10 = log x 10 = 10^x

OpenStudy (anonymous):

and x=1 ?

OpenStudy (anonymous):

the logarithm of y is defined as the exponent that the base has to be taken to equal y

OpenStudy (anonymous):

so if y = log2 100 then 100 = 2^y

OpenStudy (anonymous):

seems easy now ! :)